find the exact value of sin2x given cosx=-12/13 and x is in quadrant 3
Answer by lwsshak3(11628)

find the exact value of sin2x given cosx=-12/13 and x is in quadrant 3
**
in quadrant 3
cosx=-12/13
sin^2x=1-cos^2x=1-144/169=169/169-144/169=25/169
sinx=√(25/69)=-5/13
sin 2x=2sinx cosx=2*-5/13*-12/13=120/169